Calculate the mass percent, % (m/m), for the solute in each of the following solutions.
Answer #1

A)

mass(KCl)= 25 g


mass(solvent)= 125 g

mass of solution = 25.0+125.0 = 150.0 g

use:
mass% = mass of KCl *100 / mass of solution
= 25*100/150
= 16.67 %
Answer: 17 %

B)
mass(sugar)= 15 g


mass(total)= 300 g


use:
mass% = mass of sugar *100 / total mass
= 15*100/300
= 5.0 %
Answer: 5.0 %

c)

mass(CaCl2)= 9.0 g


mass(total)= 60.0 g


use:
mass% = mass of CaCl2 *100 / total mass
= 9.0*100/60
= 15 %
Answer: 15 %